Write the standard form of the line that contains a slope of 1/2 and y -intercept of 3. Include your work in your final answer. Type your answer in the box provided or use the upload option to submit your solution.
1
Expert's answer
2020-08-30T18:00:57-0400

Consider the standard line formula: A*x+B*y+C=0, where A, B, and C are constants.Transform the equation with respect to y. we Get y=-(A/B)*x-C/B. Let -A/B=k and -C/B=b, where k is the constant showing the slope of the line, and b is the y-intercept. Thus we get the equation of the line y=0.5*x+3. Let's bring this equation to the standard form. We get -x+2*y-6=0.


The answer to this question is: - x+2*y-6=0